# Constructing Safety Cases for AI Systems: A Reusable Template Framework

### v1

@@ Abstract
Safety cases, structured arguments that a system is acceptably safe, are becoming central to the governance of AI systems. Yet, traditional safety-case practices from aviation or nuclear engineering rely on well-specified system boundaries, stable architectures, and known failure modes. Modern AI systems such as generative and agentic AI are the opposite. Their capabilities emerge unpredictably from low-level training objectives, their behaviour varies with prompts, and their risk profiles shift through fine-tuning, scaffolding, or deployment context. This study examines how safety cases are currently constructed for AI systems and why classical approaches fail to capture these dynamics. It then proposes a framework of reusable safety-case templates, each following a predefined structure of claims, arguments, and evidence tailored for AI systems. The framework introduces comprehensive taxonomies for AI-specific claim types (assertion-based, constrained-based, capability-based), argument types (demonstrative, comparative, causal/explanatory, risk-based, and normative), and evidence families (empirical, mechanistic, comparative, expert-driven, formal methods, operational/field data, and model-based). Each template is illustrated through end-to-end patterns addressing distinctive challenges such as evaluation without ground truth, dynamic model updates, and threshold-based risk decisions. The result is a systematic, composable, and reusable approach to constructing and maintaining safety cases that are credible, auditable, and adaptive to the evolving behaviour of generative and frontier AI systems.
@@ 1 Introduction
Safety cases have become a central assurance artefact across safety-critical engineering, structured arguments supported by evidence, intended to demonstrate that a system is acceptably safe for a specific purpose and operating context [kelly1996systematic, bloomfield2009safety]. However, artificial-intelligence (AI) systems including frontier, generative and agentic AI systems fundamentally differ from the deterministic and specification-driven systems for which safety cases were originally designed. Their capabilities are not engineered line-by-line but discovered after training. Their risks evolve through interactions, fine-tuning, and context. In addition, their evaluation often proceeds without fixed ground truth. Consequently, the logic of assurance for AI should be discovery-driven, continually updated, and capable of integrating empirical, statistical, and abductive reasoning alongside classical deductive forms.
@@
The value of a safety case lies not only in regulatory compliance or as a governance tool, but as an epistemic discipline that stabilises ethical reasoning under uncertainty. Discussions of AI safety and ethics often operate largely in a pre-harm, anticipatory mode [Konigs2025Negativity], where severity is inferred rather than observed. While this is necessary, such analysis is often conducted under conditions of limited empirical grounding, making it difficult to specify risk severity, compare against existing systems, or define acceptable risk thresholds. The safety-case approach is well suited to addressing these challenges. By grounding ethical and safety concerns in explicit claims, evidence, and well-defined comparators or baselines, AI safety cases support consistent risk assessment, review, and justification across different systems and contexts.
@@
While early work has begun to adapt safety-case thinking to frontier AI [buhl2024safety, carlan2024dynamic, clymer2025example], the field still lacks a coherent, reusable structure that practitioners can apply across model types, lifecycle stages, and regulatory regimes. Existing examples tend to focus on isolated instances such as dynamic updating, misuse safeguards, or autonomous-system templates, without an overarching synthesis that ties together claim formulation, argument structure, and evidence design. Moreover, current practices rarely capture the iterative process by which AI developers discover new model capabilities and failure modes through evaluation, stress testing, and adversarial probing [wiggins2022opportunities]. These discoveries continually reshape the scope of assurance itself.
@@
This study develops a systematic framework of reusable safety-case templates for AI systems, addressing the limitations of existing assurance approaches. The key contributions are fourfold.
@@
  • Characterisation of modern AI safety cases. The study first analyses how safety cases are currently constructed for AI systems, contrasting them with classical engineering approaches and identifying distinctive features such as capability discovery, absence of ground truth, continuous evolution, and threshold-based decision-making.
@@
  • Taxonomies for reusable safety-case design. It introduces AI-specific taxonomies for Claim types including assertion-based, constrained-based, capability-based claims, Argument types including demonstrative, comparative, causal/explanatory, risk-based, and normative arguments, and Evidence families such as empirical, mechanistic, comparative, expert-driven, formal methods, model-based and operational/field data. Each linked through explicit reasoning logics (deductive, inductive, abductive, statistical, and analogical).
@@
  • Integration with dynamic assurance. Building on emerging ideas such as Dynamic Safety-Case Management Systems (DSCMS) and Checkable Safety Arguments [carlan2024dynamic], the framework embeds these templates within a continuous evaluation pipeline, linking safety claims to live metrics and governance artefacts.
These contributions establish a composable, auditable, and reusable foundation for constructing safety cases that remain credible under the uncertainty, discovery, and rapid change inherent to modern and emerging AI systems. To illustrate the practical implications of this study, we examine a real-world case study involving an AI-based tender evaluation system in government. This case study demonstrates how the proposed framework can be applied in practice, showing its potential to enhance governance and assurance processes. It also provides concrete insights that inform decision-making and guide the development of effective, responsible AI solutions.
@@
Section 2 provides background information about AI safety case, including in-depth analysis of how AI safety cases diverge from traditional engineering examples, highlighting the distinctive challenges of discovery-driven evaluation and evolving system boundaries. We then present the methodology of this study in Section 3. Section 4 introduces the reusable safety-case template framework, detailing the proposed taxonomies of claim, argument, and evidence (CAE) types. Section 5 presents end-to-end patterns illustrating how these templates address characteristic AI assurance problems such as marginal risk without ground truth and continuous model evolution. Section 6 provides a case study to show how the framework/patterns can be applied in the real-world context. In Section 7 we discuss current limitations of this study and future work. We then conclude this study in Section 8.
@@ 2 Background
In this section we introduce core AI safety case concepts including the Claims–Arguments–Evidence (CAE) framework, taxonomy, template, and pattern, presents an overview of the AI safety ecosystem pipeline involving different stakeholders, and examine the transition from traditional engineering safety cases to AI-specific approaches.
@@ 2.1 AI safety case concepts
Claims–Arguments–Evidence (CAE). Recent works build safety cases using the CAE structure[bloomfield2021templates, goemans2024template]. For example, Goemans et al. [goemans2024template] propose a safety-case template for cyber inability that explicitly uses CAE to make safety arguments coherent and explicit. The CAE framework (see Figure1) [caeframework2024] emphasizes:
  • Claim: A claim is a true/false statement about a property of a particular object. A claim is just what you might consider it to be from common usage of the term; an idea that someone is trying to convince someone else is true. An example claim could be what you are asserting (e.g., "the model cannot perform X harmful capability").
@@
  • Evidence: Evidence is an artefact that establishes facts that can be trusted and lead directly to a claim. E.g., data, test results, evaluations or simulation results that support the argument.
@@
Safety Case Taxonomy. A taxonomy denotes a categorical classification scheme for safety case elements, organizing the different types of claims, argument structures, and evidence categories that may appear in an AI safety case. Research on safety-case taxonomies spans foundational argument structures, frequent evidence types, and emerging frontier-AI–specific argument types. Kelly [kelly1996systematic] classifies reusable argument structures into top-down, bottom-up, and general construction patterns, introducing formal notions of claim decomposition and argument strategies. Nair et al. [nair2014evidence] offer a taxonomy of evidence types for safety certification, such as Hazards Cause Specification and Testing Results. Buhl et al. [buhl2024safety] identify four core safety-case components and introduce frontier-AI specific argument types, such as directly risk-based arguments and inability arguments, reflecting the shift from traditional failure-based reasoning to arguments about model behaviour and emergent capabilities.
@@
However, existing taxonomies  [clymer2024justifysafety, S21-affirmative-safety, S116-evaluation-ai-scheming] face several limitations that hinder their application to contemporary AI systems. First, the absence of standardized taxonomies leads to inconsistent terminology and incomparable safety claims across organizations, what one team calls a "robustness claim" may differ substantially from another’s interpretation. Second, existing taxonomies focus on isolated aspects (patterns, argument, evidence types, or high-level components) rather than providing an integrated view of the complete CAE framework. Third, prior work has not adequately addressed AI-specific challenges such as emerging capability discovery, continues evolution, and context dependent behaviours that distinguish AI systems from traditional safety-critical systems. We introduce our taxonomy to address critical gaps in current AI safety practices. Building on review of existing works, our CAE taxonomy serves as a systematic framework for analyzing the primary AI safety case papers (Appendix B) we studied, enabling us to identify which claim types, argument structures, and evidence categories are most prevalent in current practice.
@@
Safety Case Templates. A template denotes a reusable structured blueprint for constructing a concrete safety case in a particular context or domain. While the taxonomy provides the classification of element types, a template specifies how selected instances of these elements are arranged and link or stitch multiple CAE chains to form a coherent assurance argument.
@@
Previous work on safety case template originates from safety-critical domains such as autonomous vehicles. One of the influential work is the comprehensive set of templates proposed by Bloomfield et al. [bloomfield2021templates] outline CAE-based template for autonomous systems, including requirements templates, hazard-analysis templates, safety-monitor templates, and ML-sensor templates. Recently, template-based approaches have extended into frontier AI, Goemans et al. [goemans2024template] has developed cyber inability safety case template, which integrates model capability evaluation as a first-class component, a methodological shift required by the opacity of modern AI systems, where behavioural evidence must be empirically derived rather than analytically proven.
@@
However, existing template work  [bloomfield2021templates, goemans2024template, S34-assurance-template-principles] has limited coverage of diverse claim, argument and evidence categories. We present templates for each CAE element: Claim are structured with the top level claims to sub-claims, which claim key objectives that must be addressed to demonstrate a AI system’s safety in a specific operational context. Argument providing diverse break down types that capture the justification and reasoning logic for accepting the claim based on the evidence. Evidence presenting the data and facts that substantiate the arguments, with diverse evidence types and quality criteria to ensure robustness, traceability, and sufficiency of the safety case. Our templates provide a structured lens for analyzing existing AI safety cases and guide practitioners in systematic safety case construction.
@@
Safety Case Patterns. A pattern denotes a reusable argument fragment or reasoning structure that captures a proven approach to addressing specific safety problems. Unlike templates, which provide complete structural blueprints, patterns are modular components that can be instantiated and combined across multiple contexts, embodying recurring reasoning strategies such as red-team/blue-team evaluation or defense-in-depth arguments. Research on safety case patterns has evolved from traditional systems to AI applications. Kelly et al. [KellyM97Patterns] and Alexander et al. [Alexander2007Patterns] introduce early pattern-based methods for structuring safety arguments and supporting cross-project reuse. With the rise of ML, Wozniak et al. [WozniakCAP20PatternML] propose patterns covering data quality, model training, and runtime monitoring, while Kaur et al. [KaurICSL20PatternsDNN] extend these to deep learning with robustness testing and neural-network evaluation. Porter et al. [S11-ethics-assurance-pattern] propose principles-based patterns that frame ethical principles (fairness, transparency) as top-level claims, demonstrating broader assurance objectives beyond technical safety.
@@
We have found that existing patterns  [KellyM97Patterns, Alexander2007Patterns, WozniakCAP20PatternML, KaurICSL20PatternsDNN] remain focused on specific traditional software or ML system challenges without coverage of frontier AI risks such as emerging capability, absence of ground truth, continuous evolution behaviours, and lack explicit connections to the broader CAE framework. We introduce structured four AI safety case patterns (Section 5) to address emerging AI risks challenges.
@@ 2.2 Overview of AI safety cases ecosystem pipeline
To situate AI safety cases within real operational contexts, Figure 3 presents an overview of the pipeline that supports how AI safety cases are created, assured, and governed by different stakeholders in practice. In real deployments, safety cases are not static documents produced by developers alone; they form a shared artefact that connects multiple organisations with distinct responsibilities, including AI system developers and owners, government regulators, and independent oversight bodies. Each of these actors relies on the safety case for different purposes: developers to demonstrate risk-informed design and testing, regulators to verify compliance and defensibility before approval, and audit organisations to maintain long-term accountability across versions.
Figure 3 therefore maps coordinated components needs into a coherent pipeline centred on the evolving AI safety case, which is continuously updated with inputs from risk assessments, testing, evaluation, and operational monitoring [NOPSEMA2025safetycase, dong2024agentops2]. These three coordinated components, the AI Safety Case Builder, AI Safety Case Validator, and AI Safety Case Registry, support this lifecycle by generating initial safety cases, verifying their completeness and compliance, and preserving validated versions for ongoing oversight.
  • AI Safety Case Builder/plugins: help AI developers and system owners construct structured, versioned safety cases across the AI lifecycle. At design time, the builder imports outputs from risk assessments [S99-frontier-ai] to establish the structure of claims and required evidence, and later integrates results from testing, evaluation, and monitoring. Each design or testing iteration automatically updates the safety case, maintaining traceability between identified risks, implemented controls, and empirical results that demonstrate AI system safety and reliability.
  • AI Safety Case Validator: used by government assurance bodies or accredited third parties to examine the integrity, completeness, and consistency of submitted safety cases. The Validator could check alignment with standards such as the Australian AI Safety Standard [VAISS_V1] or ISO 42001 [ISO42001_2023], verify evidence authenticity, and assess whether controls remain effective over time. It ensures that the safety case is auditable, transparent, and defensible before approval or deployment in critical contexts.
  • AI Safety Case Registry: act as a trusted repository for storing and sharing verified safety cases across agencies and sectors. It would enable oversight organisations[fairlyAI2025trust] to compare systems, trace changes, and support continuous audit readiness.
The Builder, Validator, and Registry would form a pipeline for generating, verifying, and governing AI safety cases, making AI assurance a structured, evidence-based, and collaborative process among different stakeholders, e.g., developers, regulators and auditors.

### v3

# A Structured Approach to Safety Case Construction for AI Systems

@@ Abstract
Safety cases, structured arguments that a system is acceptably safe, are becoming central to the governance of AI systems. Yet, traditional safety-case practices from aviation or nuclear engineering rely on well-specified system boundaries, stable architectures, and known failure modes. Modern AI systems, such as generative and agentic AI, are the opposite. Their capabilities emerge unpredictably from low-level training objectives, their behaviour varies with prompts, and their risk profiles shift through fine-tuning, scaffolding, or deployment context. This study examines how safety cases are currently constructed for AI systems and why classical approaches fail to capture these dynamics. This study introduces comprehensive taxonomies for AI-specific claim types (assertion-based, constraint-based, capability-based), argument types (demonstrative, comparative, causal/explanatory, risk-based, and normative), and evidence families (empirical, mechanistic, comparative, expert-driven, formal methods, operational/field data, and model-based). It then proposes a reusable safety-case template, each of which follows a predefined structure of claims, arguments, and evidence tailored for AI systems. Each template is illustrated by end-to-end patterns that address distinctive challenges, such as evaluation without ground truth, dynamic model updates, and threshold-based risk decisions. The result is a systematic, composable, and reusable approach to constructing and maintaining safety cases that are credible, auditable, and adaptive to the evolving behaviour of generative and frontier AI systems.
@@ 1 Introduction
Safety cases have become a central assurance artefact across safety-critical engineering, structured arguments supported by evidence, intended to demonstrate that a system is acceptably safe for a specific purpose and operating context [kelly1999arguing, bloomfield2009safety]. However, artificial intelligence (AI) systems, including frontier, generative, and agentic AI systems, fundamentally differ from the deterministic, specification-driven systems for which safety cases were originally designed. Their capabilities are not engineered line by line but discovered after training. Their risks evolve through interactions, fine-tuning, and context. In addition, their evaluation often proceeds without a fixed ground truth. Consequently, the logic of assurance for AI should be discovery-driven, continually updated, and capable of integrating empirical, statistical, and abductive reasoning alongside classical deductive forms.
@@
The value of a safety case lies not only in regulatory compliance or as a governance tool, but also in being an epistemic discipline that stabilises ethical reasoning under uncertainty. Discussions of AI safety and ethics often operate largely in a pre-harm, anticipatory mode [Konigs2025Negativity], where severity is inferred rather than observed. While this is necessary, such analysis is often conducted with limited empirical grounding, making it difficult to specify risk severity, compare it with existing systems, or define acceptable risk thresholds. The safety-case approach is well-suited to addressing these challenges. By grounding ethical and safety concerns in explicit claims, evidence, and well-defined comparators or baselines, AI safety cases support consistent risk assessment, review, and justification across different systems and contexts.
@@
While early work has begun to adapt safety-case thinking to frontier AI [buhl2024safety, carlan2024dynamic, clymer2025example], the field still lacks a coherent, reusable structure that practitioners can apply across model types, lifecycle stages, and regulatory regimes. Existing examples tend to focus on isolated instances such as dynamic updating, misuse safeguards, or autonomous-system templates, without an overarching synthesis that ties together claim formulation, argument structure, and evidence design. Moreover, current practices rarely capture the iterative process by which AI developers discover new model capabilities and failure modes through evaluation, stress testing, and adversarial probing [bommasani2021opportunities]. These discoveries continually reshape the scope of assurance itself.
@@
This study develops a comprehensive taxonomy of AI safety cases and reusable safety-case templates for AI systems, addressing the limitations of existing assurance approaches. The key contributions are fourfold.
@@
  • Characterisation of modern AI safety cases. The study first analyses how safety cases are currently constructed for AI systems, contrasting them with classical engineering approaches and identifying distinctive features, including capability discovery, the absence of ground truth, continuous evolution, and threshold-based decision-making.
@@
  • Taxonomies for reusable safety-case design. It introduces AI-specific taxonomies for Claim types, including assertion-based, constrained-based, capability-based claims, Argument types including demonstrative, comparative, causal/explanatory, risk-based, and normative arguments, and Evidence families such as empirical, mechanistic, comparative, expert-driven, formal methods, model-based, and operational/field data. Each is linked through explicit reasoning logics (deductive, inductive, abductive, statistical, and analogical).
@@
  • Integration with dynamic assurance. Building on emerging ideas such as Dynamic Safety-Case Management Systems (DSCMS) and Checkable Safety Arguments [carlan2024dynamic], the study embeds these templates within a continuous evaluation pipeline, linking safety claims to live metrics and governance artefacts.
Existing AI safety case approaches typically provide domain-specific instantiations (e.g., cyber inability templates), governance-oriented guidance for frontier AI, or concepts for dynamic safety management. However, they focus on specific assurance elements and do not jointly provide an integrated CAE taxonomy spanning claims, arguments, and evidence, a reusable template meta-model, and a composable pattern library that can be instantiated across recurring AI assurance scenarios. The contributions of this study establish a composable, auditable, and reusable foundation for constructing safety cases that remain credible under the uncertainty, discovery, and rapid change inherent to modern and emerging AI systems. A detailed comparison with representative approaches is provided in Table 1 (Section 2).
To illustrate the practical implications of this study, we examine a real-world case study involving an AI-based tender evaluation system in government. This case study demonstrates the practical application of the proposed approach, highlighting its potential to enhance governance and assurance processes. It also provides concrete insights that inform decision-making and guide the development of effective, responsible AI solutions.
@@
Section 2 provides background on AI safety cases, including an in-depth analysis of how they diverge from traditional engineering examples, highlighting the distinctive challenges of discovery-driven evaluation and evolving system boundaries. We then present the methodology of this study in Section 3. Section 4 introduces the reusable safety-case templates and details the proposed taxonomies of claim, argument, and evidence (CAE) types. Section 5 presents end-to-end patterns that illustrate how these templates address characteristic AI assurance problems, such as marginal risk in the absence of ground truth and continuous model evolution. Section 6 provides a case study to show how the study can be applied in the real-world context. In Section 7, we discuss current limitations of this study and future work. We then conclude this study in Section 8.
@@ 2 Background
In this section, we introduce core AI safety case concepts, including the Claims–Arguments–Evidence (CAE) approach, taxonomy, template, and pattern, present an overview of the AI safety ecosystem pipeline involving different stakeholders, and examine the transition from traditional engineering safety cases to AI-specific approaches.
@@ 2.1 AI safety case concepts
Claims–Arguments–Evidence (CAE). Recent works have developed safety cases using the CAE structure [bloomfield2021templates, goemans2024template]. For example, Goemans et al. [goemans2024template] propose a safety-case template for cyber inability that explicitly uses CAE to make safety arguments coherent and explicit. The CAE structure (see Figure 1[caeframework2024] emphasizes:
  • Claim: A claim is a true/false statement about a property of a particular object. A claim is just what you might consider to be from common usage of the term; an idea that someone is trying to convince someone else is true. An example claim could be what you are asserting (e.g., "the model cannot perform X harmful capability").
@@
  • Evidence: Evidence is an artefact that establishes facts that can be trusted and lead directly to a claim. E.g., data, test results, evaluations, or simulation results that support the argument.
@@
Safety Case Taxonomy. A taxonomy denotes a categorical classification scheme for safety case elements, organizing the different types of claims, argument structures, and evidence categories that may appear in an AI safety case. Research on safety-case taxonomies spans foundational argument structures, frequent evidence types, and emerging frontier-AI–specific argument types. Kelly [kelly1999arguing] classifies reusable argument structures into top-down, bottom-up, and general construction patterns, introducing formal notions of claim decomposition and argument strategies. Nair et al. [nair2014evidence] provide a taxonomy of evidence types for safety certification, including Hazards Cause Specification and Testing Results. Buhl et al. [buhl2024safety] identify four core safety-case components and introduce frontier-AI specific argument types, such as directly risk-based arguments and inability arguments, reflecting the shift from traditional failure-based reasoning to arguments about model behaviour and emergent capabilities.
@@
However, existing taxonomies  [clymer2024justifysafety, wasil2024affirmativesafety, balesni2024evaluations] face several limitations that hinder their application to contemporary AI systems. First, the absence of standardized taxonomies leads to inconsistent terminology and incomparable safety claims across organizations; what one team calls a "robustness claim" may differ substantially from another’s interpretation. Second, existing taxonomies focus on isolated aspects (patterns, argument, evidence types, or high-level components) rather than providing an integrated view of the complete CAE structure. Third, prior work has not adequately addressed AI-specific challenges, including the emergence of capabilities, continuous evolution, and context-dependent behaviours that distinguish AI systems from traditional safety-critical systems. We introduce our taxonomy to address critical gaps in current AI safety practices. Building on a review of existing work, our CAE taxonomy provides a systematic basis for analyzing the primary AI safety case papers (Appendix B) we studied, enabling us to identify which claim types, argument structures, and evidence categories are most prevalent in current practice.
@@
Safety Case Templates. A template denotes a reusable structured blueprint for constructing a concrete safety case in a particular context or domain. While the taxonomy classifies element types, a template specifies how selected instances of these elements are arranged and how multiple CAE chains are linked or stitched to form a coherent assurance argument.
@@
Previous work on safety case templates originates from safety-critical domains such as autonomous vehicles. One influential work is the comprehensive set of templates proposed by Bloomfield et al. [bloomfield2021templates], which outlines CAE-based templates for autonomous systems, including requirements templates, hazard-analysis templates, safety-monitor templates, and ML-sensor templates. Recently, template-based approaches have extended into frontier AI. Goemans et al. [goemans2024template] have developed a cyber-inability safety case template that integrates model capability evaluation as a first-class component, a methodological shift required by the opacity of modern AI systems, where behavioural evidence must be empirically derived rather than analytically proven.
@@
However, existing template work  [bloomfield2021templates, goemans2024template, thomas2021template] has limited coverage of diverse claim, argument, and evidence categories. We present templates for each CAE element: Claim is structured from top-level claims to sub-claims, each of which specifies the key objectives that must be addressed to demonstrate an AI system’s safety in a specific operational context. Argument provides diverse breakdown types that capture the justification and reasoning for accepting the claim given the evidence. Evidence presenting the data and facts that substantiate the arguments, with diverse evidence types and quality criteria to ensure robustness, traceability, and sufficiency of the safety case. Our templates provide a structured lens for analyzing existing AI safety cases and guide practitioners in systematically constructing safety cases.
@@
Safety Case Patterns. A pattern denotes a reusable argument fragment or reasoning structure that captures a proven approach to addressing specific safety problems. Unlike templates, which provide complete structural blueprints, patterns are modular components that can be instantiated and combined across multiple contexts, embodying recurring reasoning strategies such as red-team/blue-team evaluation or defense-in-depth arguments. Research on safety case patterns has evolved from traditional systems to AI applications. Kelly et al. [KellyM97Patterns] and Alexander et al. [Alexander2007Patterns] introduce early pattern-based methods for structuring safety arguments and supporting cross-project reuse. With the rise of ML, Wozniak et al. [WozniakCAP20PatternML] propose patterns covering data quality, model training, and runtime monitoring, while Kaur et al. [KaurICSL20PatternsDNN] extend these to deep learning with robustness testing and neural-network evaluation. Porter et al. [porter2024principles] propose principles-based patterns that frame ethical principles (fairness, transparency) as top-level claims, demonstrating broader assurance objectives beyond technical safety.
@@
We have found that existing patterns  [KellyM97Patterns, Alexander2007Patterns, WozniakCAP20PatternML, KaurICSL20PatternsDNN] remain focused on specific traditional software or ML system challenges without coverage of frontier AI risks such as emerging capabilities, the absence of ground truth, continuous evolution behaviours, and a lack of explicit connections to the broader CAE framework. We introduce four structured AI safety case patterns (Section  5) to address emerging AI risk challenges.
@@ 2.2 Overview of AI safety cases ecosystem pipeline
To situate AI safety cases within real operational contexts, Figure 3 presents an overview of the pipeline that supports how AI safety cases are created, assured, and governed by different stakeholders in practice. In real deployments, safety cases are not static documents produced by developers alone; they form a shared artefact that connects multiple organisations with distinct responsibilities, including AI system developers and owners, government regulators, and independent oversight bodies. Each of these actors rely on the safety case for different purposes: developers to demonstrate risk-informed design and testing, and have regulators verify compliance and defensibility before approval, and audit organisations to maintain long-term accountability across versions.
Figure 3 therefore maps coordinated components’ needs into a coherent pipeline centred on the evolving AI safety case, which is continuously updated with inputs from risk assessments, testing, evaluation, and operational monitoring [NOPSEMA2025safetycase, dong2024agentops2]. These three coordinated components, the AI Safety Case Builder, AI Safety Case Validator, and AI Safety Case Registry, support this lifecycle by generating initial safety cases, verifying their completeness and compliance, and preserving validated versions for ongoing oversight.
  • AI Safety Case Builder/plugins: help AI developers and system owners construct structured, versioned safety cases across the AI lifecycle. At design time, the builder imports outputs from risk assessments [buhl2024safety] to establish the structure of claims and required evidence, and later integrates results from testing, evaluation, and monitoring. Each design or testing iteration automatically updates the safety case, maintaining traceability between identified risks, implemented controls, and empirical results that demonstrate AI system safety and reliability.
  • AI Safety Case Validator: used by government assurance bodies or accredited third parties to examine the integrity, completeness, and consistency of submitted safety cases. The Validator could check alignment with standards such as the Australian AI Safety Standard [VAISS_V1] or ISO 42001 [ISO42001_2023], verify the authenticity of evidence, and assess whether controls remain effective over time. It ensures that the safety case is auditable, transparent, and defensible before approval or deployment in critical contexts.
  • AI Safety Case Registry: acts as a trusted repository for storing and sharing verified safety cases across agencies and sectors. It would enable oversight organisations [fairlyAI2025trust] to compare systems, trace changes, and support continuous audit readiness.
The Builder, Validator, and Registry would form a pipeline for generating, verifying, and governing AI safety cases, making AI assurance a structured, evidence-based, and collaborative process among different stakeholders, e.g., developers, regulators, and auditors.